How to solve relative motion problems,and at which time we should wear which formula?
 

Grade:11

2 Answers

Arun
5 years ago
Dear Ashish
 
Please ask specific question. I can make you understand the cincept with the help of question if possible.
 
 
Regards
 
Arun
Yogansh phunni
5 years ago
When time is not given and Q. Is asking for displacement try V^2-U^2=2AS
When time is given then use S=it+1/2at^2
When Q. is asking for velocity then v=u+at where v is final velocity u is initial velocity and a is acceleration t is the time taken to complete the motion